OpenStudy (anonymous):

whAT could be a factor of n(n+1) if n is a positive integer less than 3

OpenStudy (campbell_st):

what about 1... its a factor of every number

OpenStudy (anonymous):

1 is not a factor.

OpenStudy (campbell_st):

yes it is... find the factors of 3, 1 and 3 its a factor of every number prime or composite

OpenStudy (anonymous):

the choices is 3,5,8 and 9

OpenStudy (ikram002p):

if n is a positive integer less than 3 means n=1,2 so n(n+1) =1(1+1)=2 or n(n+1) =2(1+2)=2*3

OpenStudy (ikram002p):

possible factors are , 1,2,3,6

OpenStudy (anonymous):

thank you! :)